Thinking back to our Sulfur Hexafluoride reaction, how big of a vessel would we require to store only the product in liters if we start with 2 moles of Fluorine (with just enough Sulfur to completely convert all Fluorine to the product) at standard temperatures and pressure? Assume that Sulfur Hexafluoride is an ideal gas.


1
Expert's answer
2021-10-20T02:51:26-0400

+ 24F ---> 8SF 6

as we can see from the above equation

24 mol F2 can give 8 mol SF6

or 1 mol F2 can give 8/24 mol SF6

here 2 mol of F2 is available so

number of moles of SF6 formed =(8/24)*2 =16/24 =2/3 mol SF6

we know 1 mol of an ideal gas at standard pressure and temperature has 22.4 L volume

here volume required by SF6=22.4L*(2/3) =14.9 L

so answer is 14.9L